OpenStudy (anonymous):

Factor the following binomial. x^2-225

OpenStudy (math&ing001):

x²-225=x²-15² Use a²-b²=(a-b)(a+b)

OpenStudy (anonymous):

a^2 - b^2 factors to (a - b)(a + b) 225 = 15^2...so you're question is really x^2 - 15^2...